1. Ifyou are driving 110km/h along a straight road and you look to the side for 2.0 s, how far do you travel during this inattentive period?
Answer #1

V = 110km/hr = 110×5/18 m/s

t = 2s

Hence distance s = Vt = 550×2/18 = 61.12m

In km, S = 0.0612 km